OpenStudy (kittiwitti1):

"Use a linear approximation to estimate the given number" e^(0.1) Linear approximation: L(x)=f(a)-f'(a)(x-a)

OpenStudy (kittiwitti1):

\[e^{0.1}\]function \(e^{x}\); what would I set as \(a\)?

OpenStudy (agent0smith):

I guess do it like \[\large y = e^x \]\[\large \frac{ dy }{ dx } = e^x\]\[\large \Delta y = e^x \Delta x \]then you could use e^0 as your known value, then for e^0.1 Delta x would be 0.1\[\large e^{0.1} = e^0 + \Delta y\]

OpenStudy (kittiwitti1):

I think they want me to use the L(x) which I added into the original post. Sorry about that

OpenStudy (agent0smith):

Yeah that's the same thing as what i gave, just with things plugged in.

OpenStudy (agent0smith):

\[\large L(x)=f(a)-f'(a)(x-a) \]\[\large L(0.1)=e^0-e^0(0.1-0) \]

OpenStudy (kittiwitti1):

So \(a=0\)?

OpenStudy (agent0smith):

Yes

OpenStudy (kittiwitti1):

Oh, alright. Gotcha.

OpenStudy (kainui):

@kittiwitti1 Do you understand why @agent0smith 's formula is exactly the same thing as yours?

OpenStudy (kittiwitti1):

Whoops, forgot to close this question.